Raul wants to buy a new collar for each of his 4 dogs. The collars come in a choice of 7 different colors.

Step 1 of 2 :  

How many selections of collars for the 4 dogs are possible if repetitions of colors are allowed?

Step 2 of 2 :

How many selections of collars are possible if repetitions of colors are not allowed?
Answer #1

Step - 1: Number of selections possible if repetition is allowed = 7*7*7*7 = 2401

Step - 2: Number of selections possible if repetition is not allowed = 7*6*5*4 = 840